OpenStudy (anonymous):

Solve the equation by using factoring: x2 -14x + 49 =0 A. 14 B. 7 C. 2 D. no solution

OpenStudy (anonymous):

Can't you try?

OpenStudy (anonymous):

roots are 7 and 7.. nature of roots: real, rational and equal it's C.

OpenStudy (anonymous):

you mean B: 7

OpenStudy (anonymous):

i did, i suck at math.

OpenStudy (anonymous):

There's relation between 14 and 49 :)

OpenStudy (anonymous):

lol,. yeah, it's B. .

OpenStudy (anonymous):

Yep, you need to see what two numbers add up to -14 and multiply to 49 :)

OpenStudy (anonymous):

7 was right, thanks (:

OpenStudy (anonymous):

Will you know how to do them next time? x + y = -14 x * y = 49 -7 + -7 = -14 and -7 * -7 = 49 x - 7 = 0 x = 7

OpenStudy (anonymous):

yeah i think i got it (:

OpenStudy (anonymous):

Awesome